Shift report 20011106 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Nov 06 2001 - 00:14:43 EST

  • Next message: Apache: "Shift report 20011106 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Nov  6 00:14:43 2001, Shift: 16:00-24:00
    Supervisor: DO, with: 
    Log book#10 pages 97 - 99
    
    Shift Summary:
    ------------------------------------------------------------
    Last shift report from DO!! And a good one :)
    
    1.Machine status
     18:00 injection and ramping ok, but beam lost at 18:45
     due to a permit interlock (no warning about it of course).
     19:10 new injection and nice store that was dumped only
     5mn ago by a common vote from Brahms and Phenix.
    
    2.Achievements:
      Run     trig6    Setting
      5677    33.5k    MRS @ 52.5 3A, FS @ 12 1/8 B
      5678    30.3k    same setting
      5679    52.2k    MRS @ 52.5 3A, FS @ 12 0 field
      5680   131.0k    MRS @ 52.5 3A, FS @ 12 1/8 A
     ----------------
      Total  247.0k 
    
    After dump, set MRS at 90 deg., 2B
    
    3.No problems at all.
    
    That was a really cool shift!  I still haven't fully
    recovered from my night shifts. So for those I won't
    see the next 2 days, it was nice to work with you 
    during a month. Hope to see next year for another
    nice run period.
    
    Djam  
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Nov 06 2001 - 06:54:33 EST